solve the following equation 1 over t+5+3 over t-5+=-10over t to the second-25
Answer by stanbon(57346)

solve the following equation
1 over t+5+3 over t-5+=-10over t to the second-25
-----
1/(t+5) + 3/(t-5) = -10/(t^2-25)
-----
Multiply thru by t^2-25 to get:
(t-5) + 3(t+5) = -10
---
4t+10 = -10
4t = -20
t = -5
---
But -5 cannot be a solution as that
would make 2 of the denominators zero.
---
No solution.
==================
Cheers,
Stan H.